OpenStudy (anonymous):

what is the nth term of the sequence 1,-2/3,4/9,-8/27

OpenStudy (anonymous):

(2-n)(-2/3) -4+n/3

OpenStudy (anonymous):

^ that doesn't make sense... ignore that

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

its basically asking what is the formula

OpenStudy (anonymous):

I get the pattern.. the difficult part is writing the formula... yup.

OpenStudy (anonymous):

n(-2/3)+(2/3) It's almost there... uggh... this is annoying me

OpenStudy (anonymous):

they are fractions btw

OpenStudy (anonymous):

Yeah.. I figure hahah

OpenStudy (anonymous):

figured*

OpenStudy (anonymous):

lol sorry

OpenStudy (anonymous):

it's okay... I just feel like murdering this problem.. ya know if that's physically possible. :P

OpenStudy (anonymous):

i know same here...ive been trying for so long

OpenStudy (anonymous):

(-2^n-1)/(3^n-1) n=1... 1 n=2...-2/3 n=3...4/9 n=4...-8/27 OMG I FIGURED IT OUT!! :D

OpenStudy (anonymous):

\[-2^{n-1}/3^{n-1}\]

OpenStudy (campbell_st):

the common ratio is -2/3 and a = 1 the series in geometric then \[T _{n} = 1\times(-2/3)^{n-1}\]

OpenStudy (anonymous):

^ do you really the need the one there?

OpenStudy (anonymous):

either way.. both equations work!! ^.^ and now I'm really happy because I get it! :D

OpenStudy (campbell_st):

well the general term in any geometric sequence is \[T _{n} = ar^{n-1}\]

OpenStudy (anonymous):

thank you both so much, great help